# VersaTune: An Efficient Data Composition Framework for Training Multi-Capability LLMs

### v3

@@ 5. Experiments
| LLaMA-2-7B | Uniform Distribution | 15.71 | 30.72 | 23.45 | 27.57 | 33.50 | 2.71 | 9.30 | 42.89 | 5.67 | 3.44 | 20.16 | 71.40 |
@@
| LLaMA-2-13B | Uniform Distribution | 47.66 | 34.85 | 32.98 | 36.54 | 37.54 | 32.85 | 45.72 | 50.77 | 36.54 | 38.55 | 36.89 | 73.50 |
@@ 5.1. Experimental Setup
Model Details. We use LLaMA (Dubey et al., 2024; Touvron et al., 2023a, b) and Qwen (Bai et al., 2023; Yang et al., 2024) series as our pretrained language models . For scenarios aimed at fostering multi-ability, we trained and assessed models including LLaMA-2-7B, LLaMA-3-8B, LLaMA-2-13B, Qwen-1.5-7B, Qwen-2-7B, Qwen-2.5-7B, and Qwen-2.5-14B. In the context of domain expansion, the training and evaluations were performed using the LLaMA-3-8B and LLaMA-2-13B models. Regarding reference models, for the LLaMA series, we used the Sheared-LLaMA-1.3B (Xia et al., 2023) as a lightweight reference model; as for the Qwen series, we utilized Qwen-2-1.5B and Qwen-2.5-1.5B as our reference models.
@@ 5.3. Results
| 1 | Uniform Distribution of Non-Target Domains | 5.38 | 7.35 | 6.47 | 6.58 | 0.05 | 13.99 | 5.38 | 6.48 | |
@@
| 2 | Uniform Distribution of Non-Target Domains | 33.89 | 7.94 | 26.47 | 12.65 | 7.65 | 5.00 | 33.89 | 3.23 | |
@@
| 3 | Uniform Distribution of Non-Target Domains | 52.89 | 8.24 | 15.00 | 12.06 | 10.17 | 5.12 | 52.89 | 20.59 | |
@@
| 4 | Uniform Distribution of Non-Target Domains | 58.12 | 9.71 | 5.59 | 9.41 | 10.14 | 12.05 | 58.12 | 35.72 | |
@@
| 1 | Uniform Distribution of Non-Target Domains | 3.82 | 5.36 | 5.29 | 5.59 | 2.65 | 6.53 | 5.36 | 0.36 | |
@@
| 2 | Uniform Distribution of Non-Target Domains | 7.94 | 33.68 | 9.12 | 20.59 | 4.98 | 5.64 | 33.68 | 2.87 | |
@@
| 3 | Uniform Distribution of Non-Target Domains | 10.46 | 55.69 | 14.98 | 12.35 | 4.41 | 14.18 | 55.69 | 31.68 | |
@@
| 4 | Uniform Distribution of Non-Target Domains | 13.82 | 58.07 | 11.18 | 2.35 | 6.47 | 18.53 | 58.07 | 47.65 | |
@@
| 1 | Uniform Distribution of Non-Target Domains | 12.05 | 7.36 | 8.21 | 6.74 | 19.41 | 8.53 | 8.21 | 8.83 | |
@@
| 2 | Uniform Distribution of Non-Target Domains | 20.59 | 7.94 | 31.08 | 4.71 | 7.69 | 4.98 | 31.08 | 5.23 | |
@@
| 3 | Uniform Distribution of Non-Target Domains | 15.02 | 8.24 | 52.41 | 12.06 | 10.30 | 4.98 | 52.41 | 20.56 | |
@@
| 4 | Uniform Distribution of Non-Target Domains | 5.88 | 5.29 | 56.13 | 14.09 | 13.23 | 20.87 | 56.13 | 47.60 | |
@@
| 1 | Uniform Distribution of Non-Target Domains | 6.76 | 17.64 | 9.18 | 5.37 | 7.05 | 4.73 | 5.37 | 13.48 | |
@@
| 2 | Uniform Distribution of Non-Target Domains | 8.24 | 20.59 | 5.68 | 32.78 | 9.12 | 5.85 | 32.78 | 21.30 | |
@@
| 3 | Uniform Distribution of Non-Target Domains | 9.98 | 12.06 | 15.01 | 55.78 | 4.11 | 14.13 | 55.78 | 31.17 | |
@@
| 4 | Uniform Distribution of Non-Target Domains | 13.82 | 13.53 | 11.18 | 58.46 | 6.57 | 6.47 | 58.46 | 24.51 | |
@@
| 1 | Uniform Distribution of Non-Target Domains | 3.76 | 5.09 | 7.80 | 7.68 | 5.23 | 5.65 | 5.23 | 6.86 | |
@@
| 2 | Uniform Distribution of Non-Target Domains | 7.90 | 13.49 | 9.05 | 12.03 | 38.31 | 17.76 | 38.31 | 9.19 | |
@@
| 3 | Uniform Distribution of Non-Target Domains | 9.96 | 15.07 | 8.46 | 5.33 | 55.62 | 10.04 | 55.62 | 38.20 | |
@@
| 4 | Uniform Distribution of Non-Target Domains | 14.01 | 13.86 | 6.57 | 6.66 | 58.06 | 13.93 | 58.06 | 41.71 | |
| VersaTune | 5.66 | 4.42 | 10.10 | 9.95 | 59.71 | 13.57 | 59.71 | 14.96 | ||
VersaTune is efficient across different models in both scenarios. Table 4 and Figure 5 show that VersaTune consistently outperforms other baseline methods across different models in terms of domain-specific capabilities. Compared to the uniform distribution of data across domains, VersaTune enhances downstream task performances by , which further underscores the effectiveness of our data composition strategy for enhancing the model’s overall multi-domain capabilities during the supervised fine-tuning phase. Since we have not conducted domain-specific refinement for domains outside the current five specific domains, the models’ performance gains on general benchmarks are not as noticeable. For domain expansion scenarios, VersaTune has nearly maintained training efficiency while reducing the model’s loss of competencies in other domains by comparing to 100% specific domain fine-tuning, as depicted in Table 5 and Figure 6, where we averaged the experimental results from LLaMA-3-8B and LLaMA-2-13B. Three interesting phenomena are observed from the outcomes:

### v4

@@ 5. Experiments
Baichuan-2-7B | Uniform Distribution | 15.71 | 30.72 | 23.45 | 27.57 | 33.50 | 2.71 | 9.30 | 42.89 | 5.67 | 3.44 | 20.16 | 71.40 |
@@
Baichuan-2-13B | Uniform Distribution | 47.66 | 34.85 | 32.98 | 36.54 | 37.54 | 32.85 | 45.72 | 50.77 | 36.54 | 38.55 | 36.89 | 73.50 |
@@ 5.1. Experimental Setup
Model Details. We use LLaMA (Dubey et al., 2024; Touvron et al., 2023a, b), Qwen (Bai et al., 2023; Yang et al., 2024) and Baichuan (Yang et al., 2023) series as our pretrained language models . For scenarios aimed at fostering multi-ability, we trained and assessed models including Baichuan-2-7B, LLaMA-3-8B, Baichuan-2-13B, Qwen-1.5-7B, Qwen-2-7B, Qwen-2.5-7B, and Qwen-2.5-14B. For domain expansion, the training and evaluations were performed using the Baichuan-2-7B and Baichuan-2-13B models. Regarding reference models, for the LLaMA series, we used the Sheared-LLaMA-1.3B (Xia et al., 2023) as a lightweight reference model; as for Qwen series, we utilized Qwen-2-1.5B and Qwen-2.5-1.5B as our reference models.
@@ 5.3. Results
1 | Uniform Distribution of Non-Target Domains | 5.38 | 7.35 | 6.47 | 6.58 | 0.05 | 13.99 | 5.38 | 6.48 | |
@@
2 | Uniform Distribution of Non-Target Domains | 33.89 | 7.94 | 26.47 | 12.65 | 7.65 | 5.00 | 33.89 | 3.23 | |
@@
3 | Uniform Distribution of Non-Target Domains | 52.89 | 8.24 | 15.00 | 12.06 | 10.17 | 5.12 | 52.89 | 20.59 | |
@@
4 | Uniform Distribution of Non-Target Domains | 58.12 | 9.71 | 5.59 | 9.41 | 10.14 | 12.05 | 58.12 | 35.72 | |
@@
1 | Uniform Distribution of Non-Target Domains | 3.82 | 5.36 | 5.29 | 5.59 | 2.65 | 6.53 | 5.36 | 0.36 | |
@@
2 | Uniform Distribution of Non-Target Domains | 7.94 | 33.68 | 9.12 | 20.59 | 4.98 | 5.64 | 33.68 | 2.87 | |
@@
3 | Uniform Distribution of Non-Target Domains | 10.46 | 55.69 | 14.98 | 12.35 | 4.41 | 14.18 | 55.69 | 31.68 | |
@@
4 | Uniform Distribution of Non-Target Domains | 13.82 | 58.07 | 11.18 | 2.35 | 6.47 | 18.53 | 58.07 | 47.65 | |
@@
1 | Uniform Distribution of Non-Target Domains | 12.05 | 7.36 | 8.21 | 6.74 | 19.41 | 8.53 | 8.21 | 8.83 | |
@@
2 | Uniform Distribution of Non-Target Domains | 20.59 | 7.94 | 31.08 | 4.71 | 7.69 | 4.98 | 31.08 | 5.23 | |
@@
3 | Uniform Distribution of Non-Target Domains | 15.02 | 8.24 | 52.41 | 12.06 | 10.30 | 4.98 | 52.41 | 20.56 | |
@@
4 | Uniform Distribution of Non-Target Domains | 5.88 | 5.29 | 56.13 | 14.09 | 13.23 | 20.87 | 56.13 | 47.60 | |
@@
1 | Uniform Distribution of Non-Target Domains | 6.76 | 17.64 | 9.18 | 5.37 | 7.05 | 4.73 | 5.37 | 13.48 | |
@@
2 | Uniform Distribution of Non-Target Domains | 8.24 | 20.59 | 5.68 | 32.78 | 9.12 | 5.85 | 32.78 | 21.30 | |
@@
3 | Uniform Distribution of Non-Target Domains | 9.98 | 12.06 | 15.01 | 55.78 | 4.11 | 14.13 | 55.78 | 31.17 | |
@@
4 | Uniform Distribution of Non-Target Domains | 13.82 | 13.53 | 11.18 | 58.46 | 6.57 | 6.47 | 58.46 | 24.51 | |
@@
1 | Uniform Distribution of Non-Target Domains | 3.76 | 5.09 | 7.80 | 7.68 | 5.23 | 5.65 | 5.23 | 6.86 | |
@@
2 | Uniform Distribution of Non-Target Domains | 7.90 | 13.49 | 9.05 | 12.03 | 38.31 | 17.76 | 38.31 | 9.19 | |
@@
3 | Uniform Distribution of Non-Target Domains | 9.96 | 15.07 | 8.46 | 5.33 | 55.62 | 10.04 | 55.62 | 38.20 | |
@@
4 | Uniform Distribution of Non-Target Domains | 14.01 | 13.86 | 6.57 | 6.66 | 58.06 | 13.93 | 58.06 | 41.71 | |
VersaTune | 5.66 | 4.42 | 10.10 | 9.95 | 59.71 | 13.57 | 59.71 | 14.96 |
VersaTune is efficient across different models in both scenarios. Table 4 and Figure 5 show that VersaTune consistently outperforms other baseline methods across different models in terms of domain-specific capabilities. Compared to the uniform distribution of data across domains, VersaTune enhances downstream task performances by , which further underscores the effectiveness of our data composition strategy for enhancing the model’s overall multi-domain capabilities during the supervised fine-tuning phase. Since we have not conducted domain-specific refinement for domains outside the current five specific domains, the models’ performance gains on general benchmarks are not as noticeable. For domain expansion scenarios, VersaTune has nearly maintained training efficiency while reducing the model’s loss of competencies in other domains by comparing to 100% specific domain fine-tuning, as depicted in Table 5 and Figure 6, where we averaged the experimental results from Baichuan-2-7B and Baichuan-2-13B. Three interesting phenomena are observed from the outcomes: